Quick Start Company makes 12-volt car batteries. After many years of product testing, the company knows that the average life of a Quick Start battery is normally distributed, with a mean of 46.6 months and a standard deviation of 9.5 months.

(a) If Quick Start guarantees a full refund on any battery that fails within the 36-month period after purchase, what percentage of its batteries will the company expect to replace? (Round your answer to two decimal places)


(b) If Quick Start does not want to make refunds for more than 5% of its batteries under the full-refund guarantee policy, for how long should the company guarantee the batteries (to the nearest month?)
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 46.6
std deviation   =σ= 9.5000

percentage of its batteries that company expected to replace

probability = P(X<36) = P(Z<-1.12)= 13.14%

b)

for 5th percentile critical value of z= -1.64
therefore corresponding value=mean+z*std deviation= 31.0 months